Having a server with you is like having a brilliant sysadmin with good manners employed! :)
|Posted||Thu, 27 Apr 2017 02:58 AM UTC|
|Last Update||Thu, 27 Apr 2017 02:58 AM UTC (72 hours ago)|
Host server maintenance (will restart VMs).
RimuHosting has received a Xen Security Advisory (XSA) that requires us to perform updates to some of our Xen host servers. To do this we need to stop the VMs running on affected hosts, update the host servers and restart the host servers and their VMs.
These security advisories impact only some customers running VMs on some of our Xen-based host servers. They do not impact any customers with 'bare metal' dedicated servers.
Your VM will be unavailable during the update. We expect the updates to take 45 minutes (typically) to 2 hours (less commonly) per host.
Since the update requires a VM restart, now would be a good time to check your setups to ensure that your server boots up as expected. e.g. that expected services run. e.g. that SSL certificates do not require manual pass phrase entries.
You can view the scheduled upgrade times at https://rimuhosting.com/cp/hsupgrades.jsp
If a host server is listed as "Not yet scheduled" then that host does not require an update/restart.
Customers on dedicated servers can edit the scheduled time at https://rimuhosting.com/cp/vps/hsupgrade.jsp
The XSAs will be publicly released by the Xen project team on Tue May 2 UTC. We must complete this maintenance before then.
Where possible we attempt to schedule updates at times that are off peak at the host server's location.
Customers on dedicated hosts can schedule updates for a time that suits (provided it is prior to May 2). Due the logistical demands of the update we are unable to accommodate customer preferences for times on shared hosts. Customers on shared hosts can move to dedicated VM hosts should they need to control maintenance window timing.
Please subscribe to this notice to stay updated.#
|Posted||Mon, 24 Apr 2017 16:46 PM UTC|
|Last Update||Thu, 27 Apr 2017 21:44 PM UTC (53 hours ago)|
|Affected Data Center||Dallas (TX)|
We are currently experiencing some routing issues. We are escalating this to our upstream providers.
UPDATE: the datacenter staff report they are unable to replicate. And that the packet loss is way beyond their reach. Some customer are reporting improvements. We will continue to monitor this issue.
Update Tue, 25 Apr 2017 19:55 PM UTC: Level3 reports there was a known issue. Their RFO summary is below. We have not seen any issues for over 24 hours. If you are seeing anything please open a support ticket.
"Reason for Outage (RFO) Summary: The IP NOC clarified this outage was not caused by a configuration error. The link was determined to be taking heavy errors, but the root cause is currently unknown. The IP NOC removed services from the link for stability. The equipment vendor has been engaged to assist with further investigation into the root cause. Additional action will be taken as deemed necessary and available to prevent future issues."#
|Posted||Sun, 23 Apr 2017 20:01 PM UTC|
|Last Update||Sun, 23 Apr 2017 20:01 PM UTC (6 days ago)|
The host restarted itself around 1540UTC today. There was no indication in the logs as to why.
The drives and RAID are all operating normally. All of the VMs restarted normally and are responding.
We will continue to monitor it closely.#